# Architecting Scalable Trapped Ion Quantum Computers using Surface Codes

### v1

@@ 1. Introduction
Figure 1. Quantum Charge-Coupled Device (QCCD) system. A detailed view of the QCCD hardware, where ions (grey circles) serve as qubits and are confined within an electromagnetic field known as a trap. (a) The trap is structured with different types of electrodes to position ions, including dynamic electrodes (green) for time-varying signals and shim electrodes (blue) for static potentials. Transport segments (black) and junctions (orange) allow ions to move between traps. (b) The QCCD device is controlled by a classical system interfacing with Digital-to-Analog Converters (DACs), each responsible for individual electrode voltages, enabling precise ion control (Lekitsch et al., 2017). (c) We use an abstract QCCD view for this paper.
@@
Trapped ions (TI) qubits are an important platform for realising scalable quantum computers. Over a hundred academic groups are pursuing this technology (Optics and für Experimentalphysik Universität Innsbruck, 2023), and production systems have been demonstrated by IonQ, Quantinuum, Oxford Ionics and other vendors (IonQ, 2025; Quantinuum, 2025; Ionics, 2025). Small TI systems use a monolithic architecture where all qubits are housed in the same physical trap. This design is not scalable due to control challenges and poor gate fidelities (quality of gate operations), especially beyond 30 qubits (Murali et al., 2020; Leung and Brown, 2018). Instead, a modular design where ions are distributed across many small traps is seen as a path towards scalable systems. This architecture, termed the Quantum Charge-Coupled Device (QCCD) was first proposed in 2002 (Kielpinski et al., 2002) and has been demonstrated in practice by Quantinuum (Moses et al., 2023a). Figure 1 shows an example QCCD system with four traps.
@@
To achieve a practical quantum advantage over classical computing, we require $\approx 100-1000$ algorithmic qubits with an error rate of at least $10^{-9}$ (Acharya et al., 2023), which is well beyond the limits of all known qubit technologies (IBM, 2025; AI, 2025). Therefore, we require quantum error correction (QEC). Similarly to classical error correction, where bits are redundantly encoded, QEC encodes a logical qubit across multiple physical qubits, detecting and correcting errors. The surface code (Fowler and Gidney, 2018) is among the most promising candidates for QEC codes due to its compatibility with planar architectures. In this paper, we study how surface code-based logical qubits can be efficiently implemented on QCCD hardware. Although our work focuses on the surface code, our techniques and framework are more broadly applicable.
@@
For two reasons, implementing scalable surface code logical qubits in a QCCD architecture is non-trivial. First, QCCD systems offer a rich architectural design space with a range of trap capacity (number of ions per trap), communication topology (wiring between traps) and control wiring (hardware responsible for orchestrating ion movement) choices. The performance of the surface code logical qubit and its logical error rate depend heavily on the underlying device architecture. How should device architects navigate these choices for logical qubit implementation? Second, the performance of the surface code also depends on its mapping to the hardware and the routing techniques that are used to orchestrate the movement of ions. How can we optimise these mappings across various architectures and surface code parameters? Previous work and industry roadmaps either focus on noisy intermediate-scale quantum (NISQ) workloads (Murali et al., 2020) or use manual mappings (Lekitsch et al., 2017) or only pick out a few architectural choices without rigorous architectural exploration (Valentini et al., 2024).
@@
Our work performs the first systematic design space exploration for logical qubit implementation on QCCD devices. We require an efficient compilation of surface code parity-check circuits (Figure 3) onto diverse QCCD architectures to enable architecture evaluation. Only a few compilers (Sivarajah et al., 2020; Murali et al., 2020; Saki et al., 2022) support QCCD, but they are designed for NISQ applications on small QCCD hardware (Moses et al., 2023b). We developed a novel QEC and device topology-aware mapping scheme that exploits the parallelism and structure inherent in the parity check operations in the surface code to find good mapping solutions. Our compiler maps logical qubits to hardware and then implements logical qubit instructions using low-level QCCD primitives while adhering to QCCD hardware constraints. Using this compiler, we develop a toolflow for design space exploration, shown in Figure 2. This toolflow accepts a QEC code and QCCD architecture as input and then arrives at an efficient mapping, which is used alongside architectural models and logical qubit simulations to determine metrics such as cycle time, logical error rate and data rate. We use the tool to sweep the architectural design space and select optimal designs. Our contributions are as follows:
@@
  • We identify important architectural parameters for the implementation of surface codes in QCCD systems. Unlike previous works(Murali et al., 2020), we identify that a trap capacity of two ions is surprisingly ideal even though it maximises communication operations. When paired with grid connectivity and direct wiring of electrodes to DACs, we can achieve near-optimal cycle times and low logical error rates across both small and large surface code implementations, compared to higher trap capacity configurations.
@@
  • Comparing WISE(Malinowski et al., 2023), a state-of-the-art wiring method with the standard QCCD architecture, we identify a power vs. cycle time scaling bottleneck. Existing wiring methods either offer high power with fast logical clock speeds or low power but very slow speeds. For near-term demonstrations, these techniques are sufficient. However, to scale up to hundreds of logical qubits, we require a fundamental re-design of the wiring architecture considering power consumption as part of hardware-software co-design.
@@
  • Our QEC and device topology-aware compiler offers near-optimal QEC round times, outperforming existing compilers by 3.8X (Saki et al., 2022; Murali et al., 2020). Unlike existing QEC compilers for QCCD systems, our compiler can handle large surface code implementations and scale to large trap capacities.
@@ 2. Background
We focus on the surface code due to its high code threshold and ease of hardware implementation. This is because most quantum circuits for the surface code are a regular set of parity checks, where every ancilla (red) qubit is initialised, then the ancilla has a two-qubit entanglement gate with only each of its 4 neighbouring data (blue) qubits, and finally the ancilla qubit is measured (shown on the right of Figure 3). It is a well-accepted choice for TI systems (Lekitsch et al., 2017).
@@
We use a set of primitive operations that provide the quantum gates necessary to maintain a logical qubit (Gutiérrez et al., 2019). The entangling gate is a two-qubit Mølmer–Sørenson (MS) gate (t1); the implementation details are not relevant for this paper. Single-qubit gates are rotations around the x, y, and z axis on a single isolated ion (t2-t4). In addition, there are (t5) measurements of trapped-ion qubits and (t6) qubit reset. QCCD movement techniques include (t7) shuttling (moving) an ion across a transport segment connecting one trap or junction to another, (t8) splitting (moving an ion from a trap into a segment) and (t9) merging (moving an ion from a segment into a trap). An ion must be at the end of a trap in order to split (t8), which can be done by swapping ions within a trap (via 3 two-qubit gates (t1)). The final primitives are (t10) junction crossing entry and (t11) exit, whereby ions move across junctions that connect different segments. We assume that only a single ion could reside in a junction and that only a single ion could reside in a single segment at any moment (Burton et al., 2023; Walther et al., 2012; Bowler et al., 2012).
@@ 3.1. Trap Capacity
A key architectural choice for QCCD systems is trap capacity, defined as the maximum number of qubits per trap. For example, Figure 1 shows a trap with capacity 4. There are three aspects to the choice of trap capacity. First, with high capacity, inter-ion spacing reduces and makes it difficult to address individual ions in the trap with laser controllers (Murali et al., 2020), leading to poor gate fidelity. Second, with high capacity, the need for communication operations is reduced. This can improve overall circuit fidelity due to a shorter depth and the reduction in the number of noisy operations. Third, in typical trapped-ion QC implementations, the gates within the same trap are executed serially. Although parallel two-qubit gates have been demonstrated (Figgatt et al., 2019), these gate times are 6X worse than the sequential gate times we assume and the gates have been challenging to realise beyond small scales (Murali et al., 2020). To our knowledge, current QCCD platforms (IonQ, Quantinuum) do not offer parallel two-qubit gates within a trap for this reason (Chen et al., 2024). Therefore, QCCD systems with multiple small traps can execute more gates in parallel, reducing the overall execution time.
While prior works have explored the choice of trap capacity for NISQ workloads (Murali et al., 2020), the optimal trap capacity for logical qubit implementation with QCCD systems is unknown. For surface code logical qubits, there are intuitive choices for this parameter. For example, each qubit can be mapped to a separate trap. This offers the highest possible two-qubit gate fidelity at the expense of many communication operations. Similarly, non-adjacent parity checks, shown on the right of Figure 3, can each be mapped to a trap with capacity 5. This reduces communication compared to the former case. As an extreme choice, the entire logical qubit in Figure 3 can be mapped to a single trap with capacity 31 (IonQ’s systems adopt this approach (Chen et al., 2024)). As discussed, this serialises operations and kills the inherent parallelism available.
@@ 3.2. Communication Topology
To determine the optimal trap capacity, it is crucial to consider the communication topology of the QCCD device. The choice of topology determines the number of ion transport operations (t7-t11 §3) that will be required. Ions have all-to-all connectivity within a trap, while ions in different traps are connected by shuttling paths, which are implemented using segments and junctions in hardware (Figure 1). Unlike general NISQ workloads with widely varying communication requirements (Murali et al., 2020), surface code parity-check circuits have a regular local structure. As a result, ion movement operations can remain local if the communication topology between ions preserves the structure of the surface code. For example, a grid topology, where traps are interconnected by a grid network of shuttling paths and junctions (Figure 1), closely aligns with the structure of the surface code when trap capacity is minimal (Lekitsch et al., 2017). However, the performance of the grid topology is unclear when large trap capacities are used. Further, we consider two more topologies as optimistic and pessimistic cases: an all-to-all switch topology where traps are connected using an n-way junction and a linear topology where all traps are connected to their nearest neighbour on a line. The optimistic case loosely resembles the MUSIQC architecture proposed for trapped ions (Monroe et al., 2014), and the pessimistic case resembles the architecture of Quantinuum’s current H-series devices (Moses et al., 2023a).
@@ 3.3. Control System Wiring Choices
One leading alternative is the Wiring using Integrated Switching Electronics (WISE) architecture (Malinowski et al., 2023), which integrates a switch-based demultiplexing network (bottom of Figure 4). By sharing a smaller set of DACs across many electrodes, WISE scales more favourably regarding control complexity and power consumption. However, this benefit comes with a critical trade-off: only one type (t7-t11 §3) of ion movement primitive can co-occur, restricting parallelism in ion routing.
@@ 4.1. Mapping QEC Instructions to QCCD Instructions
Surface code parity-check circuits are expressed in terms of Hadamard, CNOT and measurement operations. These operations are converted into sequences of MS operations (t1) and single-qubit rotations (t2-t4) from the QCCD toolbox (§3) using known gate identities (Figgatt, 2018). This is a straightforward intermediate-representation transformation.
@@ 4.2. Mapping Qubits to Ions
In general, the balanced graph partitioning problem is NP-complete (Garey and Johnson, 1979) and has no finite factor polynomial-time approximation when partitions must be exactly equal (Andreev and Räcke, 2006). Therefore, other compilers (Sivarajah et al., 2020; Saki et al., 2022; Murali et al., 2020) that are designed for general quantum circuits are not able to efficiently cluster qubits for large code distances. Whereas, for regular grid-like graphs typical of surface codes, our compiler can approximate a balanced partition well. We use a top-down regular partitioning of the surface code topology, as depicted in Figure 6. This minimises ancilla movement between traps because qubit neighbourhoods are preserved in the map, and the surface code only contains entanglement operations between neighbouring qubits. Minor imbalances (by 1–2 qubits) can occur due to code boundary effects.
@@
In the matching problem, the edges between clusters and traps are weighted by the distance between the centre of qubit clusters in the code topology and the trap positions in the hardware topology. The problem is solved by considering all subsets of traps with cardinality equal to the number of clusters, where, for each subset, we use the Hungarian algorithm (Kuhn, 1955) to compute the minimum perfect matching in polynomial time, and the subset with the lowest total cost is selected. For general quantum circuits, the search space can be reduced to an exponential number of trap subsets by considering only contiguous subsets (no holes) whose centres lie near the centre of all traps in the hardware. To achieve polynomial-time compilation, we further prune subsets using patterns in the boundary of the surface code topology. The compiler generalises to other scalable QEC codes, since they are expected to adhere to grid-like structures compatible with the grid QCCD communication topology, making the compiler suitable for expected real-world applications.
@@ 4.3. Ion Routing Algorithm
  • Trap capacity: Each trap has a fixed maximum ion count at any time (Murali et al., 2020; Leung and Brown, 2018).
  • Junction exclusivity: Only one ion may occupy a junction at any time (Burton et al., 2023).
  • Segment exclusivity: Only one ion may occupy a shuttling segment at any time (Walther et al., 2012; Bowler et al., 2012).
@@ 5. Modelling Logical Qubits in QCCD
Table 1. Operating parameters for QCCD systems derived from  (Gutiérrez et al., 2019). The reconfiguration steps (t7–t11) do not directly cause gate infidelity; however, they introduce idling noise and increase subsequent gate error rates due to heating, quantified using the mean vibrational energy $\bar{n}$. For our analysis, we pessimistically use the upper bound values.

### v2

@@ 1. Introduction
Figure 1. Quantum Charge-Coupled Device (QCCD) system. A detailed view of the QCCD hardware, where ions (grey circles) serve as qubits and are confined within an electromagnetic field known as a trap. (a) The trap is structured with different types of electrodes to position ions, including dynamic electrodes (green) for time-varying signals and shim electrodes (blue) for static potentials. Transport segments (black) and junctions (orange) allow ions to move between traps. (b) The QCCD device is controlled by a classical system interfacing with Digital-to-Analog Converters (DACs), each responsible for individual electrode voltages, enabling precise ion control (Lekitsch_2017). (c) We use an abstract QCCD view for this paper.
@@
Trapped ions (TI) qubits are an important platform for realising scalable quantum computers. Over a hundred academic groups are pursuing this technology (QIPGroups), and production systems have been demonstrated by IonQ, Quantinuum, Oxford Ionics and other vendors (ionq; quantinuum; oxfordionics). Small TI systems use a monolithic architecture where all qubits are housed in the same physical trap. This design is not scalable due to control challenges and poor gate fidelities (quality of gate operations), especially beyond 30 qubits (murali2020architectingnoisyintermediatescaletrapped; PhysRevA.98.032318). Instead, a modular design where ions are distributed across many small traps is seen as a path towards scalable systems. This architecture, termed the Quantum Charge-Coupled Device (QCCD) was first proposed in 2002 (Kielpinski2002) and has been demonstrated in practice by Quantinuum (H2Racetrack). Figure 1 shows an example QCCD system with four traps.
@@
To achieve a practical quantum advantage over classical computing, we require $\approx 100-1000$ algorithmic qubits with an error rate of at least $10^{-9}$ (suppressing_quantum_errors_by_scaling_surface_code), which is well beyond the limits of all known qubit technologies (IBMQuantumRoadmap; GoogleQuantumAIRoadmap). Therefore, we require quantum error correction (QEC). Similarly to classical error correction, where bits are redundantly encoded, QEC encodes a logical qubit across multiple physical qubits, detecting and correcting errors. The surface code (Fowler2018LowOQ) is among the most promising candidates for QEC codes due to its compatibility with planar architectures. In this paper, we study how surface code-based logical qubits can be efficiently implemented on QCCD hardware. Although our work focuses on the surface code, our techniques and framework are more broadly applicable.
@@
For two reasons, implementing scalable surface code logical qubits in a QCCD architecture is non-trivial. First, QCCD systems offer a rich architectural design space with a range of trap capacity (number of ions per trap), communication topology (wiring between traps) and control wiring (hardware responsible for orchestrating ion movement) choices. The performance of the surface code logical qubit and its logical error rate depend heavily on the underlying device architecture. How should device architects navigate these choices for logical qubit implementation? Second, the performance of the surface code also depends on its mapping to the hardware and the routing techniques that are used to orchestrate the movement of ions. How can we optimise these mappings across various architectures and surface code parameters? Previous work and industry roadmaps either focus on noisy intermediate-scale quantum (NISQ) workloads (murali2020architectingnoisyintermediatescaletrapped) or use manual mappings (Lekitsch_2017) or only pick out a few architectural choices without rigorous architectural exploration (valentini2024demonstrationtwodimensionalconnectivityscalable).
@@
Our work performs the first systematic design space exploration for logical qubit implementation on QCCD devices. We require an efficient compilation of surface code parity-check circuits (Figure 3) onto diverse QCCD architectures to enable architecture evaluation. Only a few compilers (Sivarajah_2020; murali2020architectingnoisyintermediatescaletrapped; muzzletheshuttle) support QCCD, but they are designed for NISQ applications on small QCCD hardware (Moses_2023). We developed a novel QEC and device topology-aware mapping scheme that exploits the parallelism and structure inherent in the parity check operations in the surface code to find good mapping solutions. Our compiler maps logical qubits to hardware and then implements logical qubit instructions using low-level QCCD primitives while adhering to QCCD hardware constraints. Using this compiler, we develop a toolflow for design space exploration, shown in Figure 2. This toolflow accepts a QEC code and QCCD architecture as input and then arrives at an efficient mapping, which is used alongside architectural models and logical qubit simulations to determine metrics such as cycle time, logical error rate and data rate. We use the tool to sweep the architectural design space and select optimal designs. Our contributions are as follows:
@@
  • We identify important architectural parameters for the implementation of surface codes in QCCD systems. Unlike previous works(murali2020architectingnoisyintermediatescaletrapped), we identify that a trap capacity of two ions is surprisingly ideal even though it maximises communication operations. When paired with grid connectivity and direct wiring of electrodes to DACs, we can achieve near-optimal cycle times and low logical error rates across both small and large surface code implementations, compared to higher trap capacity configurations.
@@
  • Comparing WISE(1000qubits), a state-of-the-art wiring method with the standard QCCD architecture, we identify a power vs. cycle time scaling bottleneck. Existing wiring methods either offer high power with fast logical clock speeds or low power but very slow speeds. For near-term demonstrations, these techniques are sufficient. However, to scale up to hundreds of logical qubits, we require a fundamental re-design of the wiring architecture considering power consumption as part of hardware-software co-design.
@@
  • Our QEC and device topology-aware compiler offers near-optimal QEC round times, outperforming existing compilers by 3.8X (muzzletheshuttle; murali2020architectingnoisyintermediatescaletrapped). Unlike existing QEC compilers for QCCD systems, our compiler can handle large surface code implementations and scale to large trap capacities.
@@ 2. Background
We focus on the surface code due to its high code threshold and ease of hardware implementation. This is because most quantum circuits for the surface code are a regular set of parity checks, where every ancilla (red) qubit is initialised, then the ancilla has a two-qubit entanglement gate with only each of its 4 neighbouring data (blue) qubits, and finally the ancilla qubit is measured (shown on the right of Figure 3). It is a well-accepted choice for TI systems (Lekitsch_2017).
@@
We use a set of primitive operations that provide the quantum gates necessary to maintain a logical qubit (transversality_lattice_surgery). The entangling gate is a two-qubit Mølmer–Sørenson (MS) gate (t1); the implementation details are not relevant for this paper. Single-qubit gates are rotations around the x, y, and z axis on a single isolated ion (t2-t4). In addition, there are (t5) measurements of trapped-ion qubits and (t6) qubit reset. QCCD movement techniques include (t7) shuttling (moving) an ion across a transport segment connecting one trap or junction to another, (t8) splitting (moving an ion from a trap into a segment) and (t9) merging (moving an ion from a segment into a trap). An ion must be at the end of a trap in order to split (t8), which can be done by swapping ions within a trap (via 3 two-qubit gates (t1)). The final primitives are (t10) junction crossing entry and (t11) exit, whereby ions move across junctions that connect different segments. We assume that only a single ion could reside in a junction and that only a single ion could reside in a single segment at any moment (Burton_2023; PhysRevLett.109.080501; PhysRevLett.109.080502).
@@ 3.1. Trap Capacity
A key architectural choice for QCCD systems is trap capacity, defined as the maximum number of qubits per trap. For example, Figure 1 shows a trap with capacity 4. There are three aspects to the choice of trap capacity. First, with high capacity, inter-ion spacing reduces and makes it difficult to address individual ions in the trap with laser controllers (murali2020architectingnoisyintermediatescaletrapped), leading to poor gate fidelity. Second, with high capacity, the need for communication operations is reduced. This can improve overall circuit fidelity due to a shorter depth and the reduction in the number of noisy operations. Third, in typical trapped-ion QC implementations, the gates within the same trap are executed serially. Although parallel two-qubit gates have been demonstrated (Figgatt_2019), these gate times are 6X worse than the sequential gate times we assume and the gates have been challenging to realise beyond small scales (murali2020architectingnoisyintermediatescaletrapped). To our knowledge, current QCCD platforms (IonQ, Quantinuum) do not offer parallel two-qubit gates within a trap for this reason (Chen_2024). Therefore, QCCD systems with multiple small traps can execute more gates in parallel, reducing the overall execution time.
While prior works have explored the choice of trap capacity for NISQ workloads (murali2020architectingnoisyintermediatescaletrapped), the optimal trap capacity for logical qubit implementation with QCCD systems is unknown. For surface code logical qubits, there are intuitive choices for this parameter. For example, each qubit can be mapped to a separate trap. This offers the highest possible two-qubit gate fidelity at the expense of many communication operations. Similarly, non-adjacent parity checks, shown on the right of Figure 3, can each be mapped to a trap with capacity 5. This reduces communication compared to the former case. As an extreme choice, the entire logical qubit in Figure 3 can be mapped to a single trap with capacity 31 (IonQ’s systems adopt this approach (Chen_2024)). As discussed, this serialises operations and kills the inherent parallelism available.
@@ 3.2. Communication Topology
To determine the optimal trap capacity, it is crucial to consider the communication topology of the QCCD device. The choice of topology determines the number of ion transport operations (t7-t11 §3) that will be required. Ions have all-to-all connectivity within a trap, while ions in different traps are connected by shuttling paths, which are implemented using segments and junctions in hardware (Figure 1). Unlike general NISQ workloads with widely varying communication requirements (murali2020architectingnoisyintermediatescaletrapped), surface code parity-check circuits have a regular local structure. As a result, ion movement operations can remain local if the communication topology between ions preserves the structure of the surface code. For example, a grid topology, where traps are interconnected by a grid network of shuttling paths and junctions (Figure 1), closely aligns with the structure of the surface code when trap capacity is minimal (Lekitsch_2017). However, the performance of the grid topology is unclear when large trap capacities are used. Further, we consider two more topologies as optimistic and pessimistic cases: an all-to-all switch topology where traps are connected using an n-way junction and a linear topology where all traps are connected to their nearest neighbour on a line. The optimistic case loosely resembles the MUSIQC architecture proposed for trapped ions (PhysRevA.89.022317), and the pessimistic case resembles the architecture of Quantinuum’s current H-series devices (H2Racetrack).
@@ 3.3. Control System Wiring Choices
One leading alternative is the Wiring using Integrated Switching Electronics (WISE) architecture (1000qubits), which integrates a switch-based demultiplexing network (bottom of Figure 4). By sharing a smaller set of DACs across many electrodes, WISE scales more favourably regarding control complexity and power consumption. However, this benefit comes with a critical trade-off: only one type (t7-t11 §3) of ion movement primitive can co-occur, restricting parallelism in ion routing.
@@ 4.1. Mapping QEC Instructions to QCCD Instructions
Surface code parity-check circuits are expressed in terms of Hadamard, CNOT and measurement operations. These operations are converted into sequences of MS operations (t1) and single-qubit rotations (t2-t4) from the QCCD toolbox (§3) using known gate identities (figgatt2018building). This is a straightforward intermediate-representation transformation.
@@ 4.2. Mapping Qubits to Ions
In general, the balanced graph partitioning problem is NP-complete (garey1979computers) and has no finite factor polynomial-time approximation when partitions must be exactly equal (balanced_graph_partitioning). Therefore, other compilers (Sivarajah_2020; muzzletheshuttle; murali2020architectingnoisyintermediatescaletrapped) that are designed for general quantum circuits are not able to efficiently cluster qubits for large code distances. Whereas, for regular grid-like graphs typical of surface codes, our compiler can approximate a balanced partition well. We use a top-down regular partitioning of the surface code topology, as depicted in Figure 6. This minimises ancilla movement between traps because qubit neighbourhoods are preserved in the map, and the surface code only contains entanglement operations between neighbouring qubits. Minor imbalances (by 1–2 qubits) can occur due to code boundary effects.
@@
In the matching problem, the edges between clusters and traps are weighted by the distance between the centre of qubit clusters in the code topology and the trap positions in the hardware topology. The problem is solved by considering all subsets of traps with cardinality equal to the number of clusters, where, for each subset, we use the Hungarian algorithm (Kuhn1955Hungarian) to compute the minimum perfect matching in polynomial time, and the subset with the lowest total cost is selected. For general quantum circuits, the search space can be reduced to an exponential number of trap subsets by considering only contiguous subsets (no holes) whose centres lie near the centre of all traps in the hardware. To achieve polynomial-time compilation, we further prune subsets using patterns in the boundary of the surface code topology. The compiler generalises to other scalable QEC codes, since they are expected to adhere to grid-like structures compatible with the grid QCCD communication topology, making the compiler suitable for expected real-world applications.
@@ 4.3. Ion Routing Algorithm
  • Trap capacity: Each trap has a fixed maximum ion count at any time (murali2020architectingnoisyintermediatescaletrapped; PhysRevA.98.032318).
  • Junction exclusivity: Only one ion may occupy a junction at any time (Burton_2023).
  • Segment exclusivity: Only one ion may occupy a shuttling segment at any time (PhysRevLett.109.080501; PhysRevLett.109.080502).
@@ 5. Modelling Logical Qubits in QCCD
Table 1. Operating parameters for QCCD systems derived from  (transversality_lattice_surgery). The reconfiguration steps (t7–t11) do not directly cause gate infidelity; however, they introduce idling noise and increase subsequent gate error rates due to heating, quantified using the mean vibrational energy $\bar{n}$. For our analysis, we pessimistically use the upper bound values.